OpenStudy (anonymous):

how to solve x(xy+y+1)dy-ydx=0

OpenStudy (anonymous):

The way it's written seems to suggest that it might be exact

OpenStudy (anonymous):

it doesn't work on exact method

OpenStudy (anonymous):

What are you solving for?

OpenStudy (anonymous):

y

OpenStudy (anonymous):

Hmmm let me think for a bit...\[ x(xy+y+1)dy-ydx=x(xy+y+1)y' dx-ydx \]

OpenStudy (anonymous):

\[ x(xy+y+1)dy=ydx \\ x(xy+y+1)y'=y \\ \frac{y'}{y} = \frac{1}{x(xy+y+1)} \\ (\ln(y))'=\frac{1}{x(xy+y+1)} \\ \ln(y) = \int \frac{1}{x(xy+y+1)} dx \]